OpenStudy (anonymous):

D= [-7 2 ] Find -3D [ 4 3 ] [-1 0 ] A. [ 21 2 ] [ -12 3 ] [ 3 0 ] B. No solution C. [ 21 -6 ] [ 4 3 ] [ -1 0 ] D. [ 21 -6 ] [ -12 -9 ] [ 3 0 ]

OpenStudy (anonymous):

Each and every term in the matrix will get multiplied by (-3).. Can you solve it further???

OpenStudy (anonymous):

21 -6 -12 -9 3 0

OpenStudy (anonymous):

option D

OpenStudy (anonymous):

well said waterineyes

OpenStudy (anonymous):

Thanks @karthik_a.k.a_joker..

OpenStudy (earthcitizen):

\[-3 \times \left[\begin{matrix}-7 & 2 \\ 4 & 3 \\-1& 0 \end{matrix}\right]\]

OpenStudy (anonymous):

The answer is D, you need to times every item in the matrix by -3. From this you can determine the right answer by singling out the answer.

OpenStudy (earthcitizen):

http://www.wolframalpha.com/input/?i=-3+*+ {{-7%2C2}%2C{4%2C3}%2C{-1%2C0}}

OpenStudy (anonymous):

ya answer is D